(20) Given y = 4°-5x+7, compute the rate of change in y relative to a when x = 2
Answer #1

Given:

y=\frac{4x^{3}-5x+7}{x-1}

To calculate the rate of change in y relative to x we need to calculate dy/dx.

Using quotient rule:

\frac{dy}{dx}=\frac{(12x^{2}-5)(x-1)-(1)(4x^{3}-5x+7)}{(x-1)^{2}}

Putting x = 2 (given):

\frac{dy}{dx}=\frac{(12\times 2^{2}-5)(2-1)-(1)(4 \times 2^{3}-5(2)+7)}{(2-1)^{2}}

\frac{dy}{dx}=\frac{(43)(1)-(1)(29)}{(1)^{2}}

\frac{dy}{dx}=14

So, the rate of change in y relative to x when x = 2 is 14.